OpenStudy (anonymous):

Determine whether the statement is sometimes, always, or never true. Explain your reasoning. If A and B are realnumbers, then negative times absolute value of a plus 2 times b is negative

OpenStudy (anonymous):

always negative times the absolute value of anything is negative.

OpenStudy (anonymous):

If you mean: \[-\left| A \right| + 2*B < 0\] Then it is sometimes true, some examples: A = 1, B = 2 \[-(1) + 2*(2) = 3\] which of course is not negative, so in this case is false or A = 5, B = 1 \[-(5)+2*(1) = -3\] which is negative and in this case is true.

OpenStudy (anonymous):

the whole thing is absolute value

OpenStudy (anonymous):

In that case, mfleisch5's answer is bang on.